# Analysis of Asynchronous Protocols for Entanglement Distribution in Quantum Networks

### v1

@@ Abstract
The distribution of entanglement in quantum networks is typically approached under idealized assumptions such as perfect synchronization and centralized control, while classical communication is often neglected. However, these assumptions prove impractical in large-scale networks. In this paper, we present a pragmatic perspective by exploring two minimal asynchronous protocols: a parallel scheme generating entanglement independently at the link level, and a sequential scheme extending entanglement iteratively from one party to the other. Our analysis incorporates non-uniform repeater spacings and classical communications and accounts for quantum memory decoherence. We evaluate network performance using metrics such as entanglement bit rate, end-to-end fidelity, and secret key rate for entanglement-based quantum key distribution. Our findings suggest the sequential scheme’s superiority due to comparable performance with the parallel scheme, coupled with simpler implementation. Additionally, we propose a cutoff strategy to improve performance by discarding attempts with prolonged memory idle time, effectively eliminating low-quality entanglement links. Finally, we apply our methods to the real-world topology of SURFnet and report the performance as a function of memory coherence time.
@@ I Introduction
Given that, asynchronous approaches [24] are proposed as a scalable solution with the possibility of distributed routing [25, 26, 27, 28]. In this paper, we study two types of asynchronous approaches: Sequential and parallel as shown in Figure 1. The main difference between these protocols and synchronous ones is that elementary entanglement generation is attempted independently for each direct link; hence, there is no need for perfect arrangement so that qubits arrive at each repeater precisely at the same time or within a short time window. Unlike previous studies, we not only include the effect of noisy quantum memory in our analysis but also take into account non-uniform repeater spacing and latencies due to classical communication. To avoid further complexity however, we focus on a minimal hardware setup based on first-generation repeaters, where there is only one memory (within each repeater) per optical link connecting two quantum repeaters.
@@ II-B Noise model and assumptions
$\exp\left[-\frac{1}{\tau_{\text{coh}}}\left(T_{A}+T_{B}+\sum_{i=1}^{n+1}(t_{i,L}+t_{i,R})\right)\right],$ (8)
@@ II-E Parallel protocol
Similarly, the parallel entanglement generation protocol starts by a user (sender) demanding to establish an EPR pair with another one (receiver). In this case, all repeaters on a directed path connecting the sender to the receiver simultaneously attempt to generate link-level entanglements with their successive repeater (see Figure 1(b)), and hence the namesake parallel. Each repeater then sends a classical acknowledgement signal to the corresponding repeater from which they receive a qubit. Once both quantum memories within a repeater is successfully entangled with two other repeaters, the Bell-state measurement is performed, and the outcome is passed on to the sender. An iteration is finished when the sender receives all a Bell-state measurement outcomes. Therefore, the duration of each attempt is found by
$T=\max(\{T_{i}+\sum_{j<i}\tau_{j}|1\leq i\leq n\}),$ (14)
@@ III Evaluation
For the most part in our simulations, we set the $\mu_{i}=\mu=1$ and $F_{i}=1$ in (7) and (8) for simplicity, and focus exclusively on the impact of coherence time $\tau_{\text{coh}}$ and cutoff time $\tau_{\text{cut}}$. Furthermore, we explore the effect of $\mu_{i}=\mu$, $F_{i}$, and $p_{\text{link}}$ on the feasible regime for QKD in two different protocols.
@@ III-C Feasible region for imperfect repeaters
III-D Random repeater placement
@@
III-E Different protocols performance in a real network
@@ IV Related work
Kamin et. al. [24] analyze the ebit and secret key rates for sequential and a protocol so-called “doubling” scheme where the segments are divided into two equal halves and when both halves have finally distributed an entangled state the last swapping is attempted [40]. A few papers use Markov Chain and reinforcement learning to obtain the optimal policy for elementary link-level generation [41] or to model the repeater chain evolving states and derive optimal policies for entanglement distribution in repeater chains [42, 43]. A policy will indicate when certain actions such as swap operations should be performed at intermediate repeaters. Goodenough et. al. [44] analyze the expected fidelity with or without cut-off strategies in homogeneous swap ASAP repeater chains.

### v2

@@ Abstract
The distribution of entanglement in quantum networks is typically approached under idealized assumptions such as perfect synchronization and centralized control, while classical communication is often neglected. However, these assumptions prove impractical in large-scale networks. In this paper, we present a pragmatic perspective by exploring two minimal asynchronous protocols: a parallel scheme generating entanglement independently at the link level, and a sequential scheme extending entanglement iteratively from one party to the other. Our analysis incorporates non-uniform repeater spacings and classical communications and accounts for quantum memory decoherence. We evaluate network performance using metrics such as entanglement bit rate, end-to-end fidelity, and secret key rate for entanglement-based quantum key distribution. Our findings suggest the sequential scheme’s superiority due to comparable performance with the parallel scheme, coupled with simpler implementation. Additionally, we impose a cutoff strategy to improve performance by discarding attempts with prolonged memory idle time, effectively eliminating low-quality entanglement links. Finally, we apply our methods to the real-world topology of SURFnet and report the performance as a function of memory coherence time.
@@ I Introduction
Given that, asynchronous approaches [24] are proposed as a scalable solution with the possibility of distributed routing [25, 26, 27, 28]. In this paper, we study two types of asynchronous approaches: Sequential and parallel as shown in Figure 1. The main difference between these protocols and synchronous ones is that elementary entanglement generation is attempted independently for each direct link; hence, there is no need for perfect arrangement so that qubits arrive at each repeater precisely at the same time or within a short time window. Unlike previous studies, we not only include the effect of noisy quantum memory in our analysis but also take into account non-uniform repeater spacing and latencies due to classical communication. To avoid further complexity however, we focus on a minimal hardware setup based on first-generation repeaters, where there is only one memory (within each repeater) per optical link connecting two quantum repeaters and the entanglement sources are located within each repeater.
@@ II-B Noise model and assumptions
$\exp\left[-\frac{1}{\tau_{\text{coh}}}\left(T_{A}+T_{B}+\sum_{i=1}^{n}(t_{i,L}+t_{i,R})\right)\right],$ (8)
@@ II-E Parallel protocol
Similarly, the parallel entanglement generation protocol starts by a user (sender) demanding to establish an EPR pair with another one (receiver). In this case, all repeaters on a directed path connecting the sender to the receiver simultaneously attempt to generate link-level entanglements with their successive repeater (see Figure 1(b)), and hence the namesake parallel. Each repeater then sends a classical acknowledgement signal to the corresponding repeater from which they receive a qubit. Once both quantum memories within a repeater is successfully entangled with two other repeaters, the Bell-state measurement is performed, and the outcome is passed on to the sender. An iteration is finished when the sender receives all Bell-state measurement outcomes. Therefore, the duration of each attempt is found by
$T=\max(\{T_{i}+\sum_{j\leq i}\tau_{j}|1\leq i\leq n\}),$ (14)
@@ III Evaluation
For most of the simulations, we set the $\mu_{i}=\mu=1$ and $F_{i}=1$ in (7) and (8) for simplicity, and focus exclusively on the impact of coherence time $\tau_{\text{coh}}$ and cutoff time $\tau_{\text{cut}}$. Furthermore, we explore the effect of $\mu_{i}=\mu$, $F_{i}$, and $p_{\text{link}}$ on the feasible regime for QKD in two different protocols.
@@ III-C Feasible region for imperfect repeaters
III-D The effect of classical communication delay
In this experiment, we demonstrate how considering classical communication delay can reduce the overall SKR between the two end nodes. We examine a fixed number of repeaters while increasing the end-to-end distance ($L_{e2e}$) and plot the SKR with and without classical communication delay. In the scenario without classical communication delay, we neglect the time needed to acknowledge entanglement generation and to send the swap results to the end nodes and assume that this information can be exchanged instantaneously.
Figure 5 shows the impact of classical communication and how neglecting it gives overestimates for the SKR and fidelity. Here, the solid and dashed lines of the same color represent the same protocol with and without classical communication delay, respectively. As we go to larger distances, the reduction due to the classical communication delay increases further. For example, at $500$km neglecting classical comm causes a significant overestimation where the average performance of each protocol is off by two orders of magnitude. Similarly, the overestimated values of fidelity leads to false impression that entanglement links maintain their quantum properties (i.e., $F_{e2e}>0.5$) up to distances of $400$ km or $600$ km for the sequential and parallel schemes, respectively, while considering the classical communication delays lower these bounds by at least $100$ km.
Figure 5: SKR (a) and fidelity (b) with and without classical comm delay. The solid and dashed lines of the same color represent the same protocol with and without classical comm delay, respectively.
III-E Random repeater placement
@@
III-F Different protocols performance in a real network
@@ IV Related work
Kamin et al. [24] analyze the ebit and secret key rates for sequential and another protocol called the “doubling” scheme where the segments are divided into two equal halves and when both halves have finally distributed an entangled state the last swapping is attempted [40]. A few papers use Markov Chain and reinforcement learning to obtain the optimal policy for elementary link-level generation [41] or to model the repeater chain evolving states and derive optimal policies for entanglement distribution in repeater chains [42, 43]. A policy will indicate when certain actions such as swap operations should be performed at intermediate repeaters. Goodenough et al. [44] analyze the expected fidelity with or without cut-off strategies in homogeneous swap as soon as possible (ASAP) repeater chains.
@@ VI Acknowledgements
Appendix A Derivation details for sequential protocol
In this appendix, we provide some derivation details for the sequential protocol. We consider a linear network with $n$ repeaters placed between two end users and $n+1$ links connecting them. We start by noting the moment when each repeater memory is initialized
$t_{\text{init},k,L}$ $=\sum_{i=1}^{k}(2N_{i}-1)\tau_{i},$ (15)
$t_{\text{init},k,R}$ $=t_{\text{init},k,L}+2(N_{k+1}-1)\tau_{k+1}.$ (16)
Similarly, for the end user memories we have
$t_{\text{init},A}$ $=2(N_{1}-1)\tau_{1},$ (17)
$t_{\text{init},B}$ $=\sum_{i=1}^{n+1}(2N_{i}-1)\tau_{i}.$ (18)
As a result, the total time for the end-to-end entanglement generation is given by
$T=t_{\text{init},B}+\tau_{e2e}=\sum_{i=1}^{n+1}2\tau_{i}N_{i},$ (19)
where we added $\tau_{e2e}=\sum_{i=1}^{n+1}\tau_{i}$ to account for the classical signal transmission (acknowledgement) from $B$ to $A$.
To compute the average fidelity or QBER, we need to find memory idle times individually,
$T_{A}$ $=T-2(N_{1}-1)\tau_{1},\qquad T_{B}=\tau_{e2e},$ (20)
for the end memories, and
$t_{i,L}$ $=2N_{i+1}\tau_{i+1},\qquad t_{i,R}=2\tau_{i+1},$ (21)
for the repeater memories. To calculate the expectation value of $f_{e2e}$ in (8) we need to evaluate the expectation value of the exponential term $\boldsymbol{E}(e^{-\tau_{\text{idle}}/\tau_{\text{coh}}})$. For the end-to-end fidelity we include the end memory idle times and the exponent is simplified into
$\tau_{\text{idle}}^{(F)}$ $=T_{A}+T_{B}+\sum_{i=1}^{n}(t_{i,L}+t_{i,R})$
$=3\tau_{e2e}+4\sum_{i=2}^{n+1}N_{i}\tau_{i},$ (22)
whereas for the SKR we do not include the end memory idle times, and we obtain
$\tau_{\text{idle}}^{\text{(SKR)}}=\sum_{i=1}^{n}(t_{i,L}+t_{i,R})$ $=2\sum_{i=2}^{n+1}(N_{i}+1)\tau_{i}.$ (23)
In what follows, we discuss the average values of various quantities for both cases with and without the cutoff.
A-A No cutoff
In this case, $N_{i}$ varies over all positive integers. Hence, the average time for end-to-end entanglement is simply given by
$\overline{T}=\boldsymbol{E}(T)$ $=\sum_{i=1}^{n+1}2\tau_{i}\cdot\sum_{N_{i}=1}^{\infty}N_{i}p_{i}(1-p_{i})^{N_{i}-1}=\sum_{i=1}^{n+1}\frac{2\tau_{i}}{p_{i}}.$ (24)
where in the first equality we use (19). This is a proof for expression (13) of the main text. Next, we calculate the expectation value of the exponential term for the fidelity
$\boldsymbol{E}(e^{-\tau_{\text{idle}}^{(F)}/\tau_{\text{coh}}})=e^{-3\tau_{e2e}/\tau_{\text{coh}}}\prod_{i=2}^{n+1}\left(\frac{p_{i}e^{-4\tau_{i}/\tau_{\text{coh}}}}{1-q_{i}e^{-4\tau_{i}/\tau_{\text{coh}}}}\right),$ (25)
where $q_{i}=1-p_{i}$. A similar expression can be derived for the exponential term to calculate the SKR,
$\boldsymbol{E}(e^{-\tau_{\text{idle}}^{(\text{SKR})}/\tau_{\text{coh}}})=\prod_{i=2}^{n+1}\left(\frac{p_{i}e^{-4\tau_{i}/\tau_{\text{coh}}}}{1-q_{i}e^{-2\tau_{i}/\tau_{\text{coh}}}}\right).$ (26)
We then plug in the above expressions to (8) to compute $\overline{f_{e2e}}$ which is ultimately used in (9) and (II-C) to calculate the end-to-end entanglement fidelity and the secret key fraction, respectively.
A-B With cutoff
Unlike the case without any cutoff, here there is a chance that an iteration may not finish with an end-to-end entanglement generation (i.e., no success). We start with the one repeater case as a warm-up example and next present the general case with $n$ repeaters.
A-B1 One repeater
The protocol for two segments (one repeater) is as follows:
  • 1.
    We try the first link and repeat until we succeed. Then we go to step 2.
  • 2.
    We try the second link only $m$ times, where $m=[\frac{\tau_{\text{cut}}}{2\tau_{2}}]$, $\tau_{\text{cut}}$ is the cutoff time, and $[x]$ is the integer part of $x$. If we succeed we claim a successful event, otherwise it is a failure and we go to step 1 and start from the beginning. This is to ensure that the idle time of the left memory inside the repeater does not exceed the cutoff time.
To compute the average time, we need to enumerate successful events, how long they take, and what is their success probability. Hence, the average time is found by
$\overline{T}^{\text{(cut)}}_{2}$ $=\sum_{r=1}^{\infty}P_{2,m}Q_{2,m}^{r-1}\left(\frac{2r\tau_{1}}{p_{1}}+(r-1)\tau_{\text{cut}}+\frac{2N_{m}(p_{2})}{P_{2,m}}\tau_{2}\right)$
$=\frac{2\tau_{1}}{p_{1}P_{2,m}}+\left(\frac{1}{P_{2,m}}-1\right)\tau_{\text{cut}}+\frac{2N_{m}(p_{2})}{P_{2,m}}\tau_{2},$ (27)
where $Q_{2,m}=(1-p_{2})^{m}$, $P_{2,m}=1-Q_{2,m}$, and $N_{m}(\cdot)$ is given by
$N_{m}(p)$ $=\sum_{N=1}^{m}pq^{N-1}N=\frac{1-(1+mp)q^{m}}{p}.$ (28)
It is easy to see that as $m\to\infty$ we obtain $\overline{T}^{\text{(cut)}}_{2}\to 2(\frac{\tau_{1}}{p_{1}}+\frac{\tau_{2}}{p_{2}})$.
A-B2 $n$ repeaters
Following the idea behind the underlying ebit generation protocol (that is sequential) we can calculate the average total time progressively. We have seen how to calculate the total time for two segments. Each term in the sum in (27) represents a successful event after trying the sequential scheme with cutoff $r$-times. Hence, we calculate the average time of these events by plugging the average total time for the first link (that is $\frac{2\tau_{1}}{p_{1}}$) and add the contribution from the second link (which is subject to the cutoff policy). Similarly, we can write a recursion relation for the average time of delivering entanglement to the $k$-th repeater in terms of the average time of delivering entanglement to the $(k-1)$-th repeater as follows
$T^{\text{(cut)}}_{k}=\frac{T^{\text{(cut)}}_{k-1}}{P_{k,m_{k}}}+\left(\frac{1}{P_{k,m_{k}}}-1\right)\tau_{\text{cut}}+\frac{2N_{m_{k}}(p_{k})}{P_{k,m_{k}}}\tau_{k},$ (29)
where $Q_{k,m_{k}}=(1-p_{k})^{m_{k}}$, $P_{k,m_{k}}=1-Q_{k,m_{k}}$, and $m_{k}=[\frac{\tau_{\text{cut}}}{2\tau_{k}}]$ correspond to the $k$-th segment between the $k$-th and $(k-1)$-th repeaters. In our analysis, we numerically calculate the average end-to-end ebit generation time $T^{\text{(cut)}}_{n}$ recursively by using the above equation starting with $T^{\text{(cut)}}_{1}=\frac{2\tau_{1}}{p_{1}}$.
Next, we compute the expectation value of the exponential term for the fidelity and SKR. We note that the overall memory idle time still follows (22) and (23), but are subject to the constraints $N_{i}\leq m_{i}=[\frac{\tau_{\text{cut}}}{2\tau_{i}}]$. Hence, we truncate the sum as in $\boldsymbol{E}^{\text{(cut)}}(x_{N_{i}})=P_{i,m_{i}}^{-1}\sum_{N_{i}=1}^{m_{i}}p_{i}(1-p_{i})^{N_{i}-1}x_{N_{i}}$ where the prefactor is a normalization constant introduced earlier. As a result, we obtain
$\boldsymbol{E}^{\text{(cut)}}(e^{-\tau_{\text{idle}}^{(F)}/\tau_{\text{coh}}})$ (30)
$=e^{-3\tau_{e2e}/\tau_{\text{coh}}}\prod_{i=2}^{n+1}\left(\frac{p_{i}e^{-4\tau_{i}/\tau_{\text{coh}}}}{P_{i,m_{i}}}\frac{(1-q_{i}^{m_{i}}e^{-4m_{i}\tau_{i}/\tau_{\text{coh}}})}{(1-q_{i}e^{-4\tau_{i}/\tau_{\text{coh}}})}\right),$
for the fidelity and
$\boldsymbol{E}^{\text{(cut)}}(e^{-\tau_{\text{idle}}^{(\text{SKR})}/\tau_{\text{coh}}})$ (31)
$=\prod_{i=2}^{n+1}\left(\frac{p_{i}e^{-4\tau_{i}/\tau_{\text{coh}}}}{P_{i,m_{i}}}\frac{(1-q_{i}^{m_{i}}e^{-2m_{i}\tau_{i}/\tau_{\text{coh}}})}{(1-q_{i}e^{-2\tau_{i}/\tau_{\text{coh}}})}\right).$
for the SKR. We note that in both cases taking $m_{i}\to\infty$ we recover (25) and (26), respectively.
Appendix B Derivation details for parallel protocol
In this appendix, we derive explicit equations for the swapping schedule and memory idle time in terms of set of random variables $N_{i}$ which denote the number of attempts until we get a successful entanglement link as explained in Sec. II-B. To evaluate the average quantities, we numerically sample $N_{i}$ according to the geometric distribution $p_{i}(1-p_{i})^{N_{i}}$. Here, we start with the case of one repeater as a warm up example and present the general case afterward.
B-1 one repeater
The entanglement swapping is performed at $T_{\text{sw}}=\max((2N_{1}-1)\tau_{1},2N_{2}\tau_{2})$. The total time is then given by $T=\tau_{1}+T_{\text{sw}}$, where $\tau_{1}$ is added to account for the classical communication to send the entanglement swapping measurement outcome to the sender. The idle time for the end-node memories are found to be
$T_{A}$ $=T-2(N_{1}-1)\tau_{1},$ (32)
$T_{B}$ $=T-(2N_{2}-1)\tau_{2},$ (33)
and for network memories are given by
$t_{1,L}$ $=T_{\text{sw}}-(2N_{1}-1)\tau_{1},$ (34)
$t_{1,R}$ $=T_{\text{sw}}-2(N_{2}-1)\tau_{2}.$ (35)
Hence, the total memory idle time is
$T_{A}+T_{B}+t_{1,L}+t_{1,R}=4T-(4N_{1}-1)\tau_{1}-(4N_{2}-3)\tau_{2}.$ (36)
B-2 n repeaters
In this case, there are $n+1$ elementary links connecting the $n$ repeaters. We first note that the entanglement swapping at repeater $k$ occurs at
$T_{k}=\max((2N_{k}-1)\tau_{k},2N_{k+1}\tau_{k+1}).$ (37)
as explained in the main text. The total time is given by the last classical signal which is the measurement outcome of the entanglement swapping arriving at the sender as in (14). Furthermore, the idle time of memories in the $k$-th repeater between segments $k$ and $k+1$ is given by
$t_{k,L}$ $=T_{k}-(2N_{k}-1)\tau_{k},$ (38)
$t_{k,R}$ $=T_{k}-2(N_{k+1}-1)\tau_{k+1}.$ (39)
Hence, we have
$t_{k,L}+t_{k,R}$ $=2T_{k}-(2N_{k}-1)\tau_{k}-2(N_{k+1}-1)\tau_{k+1}$
$=|(2N_{k}-1)\tau_{k}-2N_{k+1}\tau_{k+1})|+2\tau_{k+1}.$ (40)
The idle time for the end-node memory are given by
$T_{A}$ $=T-2(N_{1}-1)\tau_{1},$ (41)
$T_{B}$ $=T-(2N_{n+1}-1)\tau_{n+1}.$ (42)
For each random iteration, we plug in the numerical values of the above expressions to (8) to compute $\overline{f_{e2e}}$ which is ultimately used in (9) and (II-C). Next, we average over them to numerically calculate the expectation values of the end-to-end entanglement fidelity and the secret key fraction, respectively.
We introduce the cutoff by imposing constraints $t_{k,s}\leq\tau_{\text{cut}}$ with $s=L,R$ on the memory idle times (38) and (39) and discard those attempts which violate the constraints.